Make New Entry,
Make Followup Entry
User name dhamilton
Log entry time 17:10:17 on November 17, 2008
Entry number 249421
This entry is a followup to: 249419
keyword=Dipole back up to 2.35 GeV/c
Heidi was here and advised that the He level reading for the dipole on
the main Hall A Tools screen is 10% lower than the reading on the the
equivalent cryo/ESR tools screen. This is leading to false alarms.
When the He level got back to around 60% following the early quenching
earlier, the dipole was ramped up to the set value of 2.35 GeV/c. The NMR
is now locked and there are no alarms.